the energy absorbed by each molecule(A2) of a substance is 4.4×10-19J and bond energy per molecule is 4.0×10-19J the kinetic energy of the molecule per atom will be:
  • 2.2×10-19J
  • 2.0×10-19J
  • 2.0×10-20J
  • 4.0×10-20J

Grade:12th Pass

3 Answers

Arun
6 years ago
That will be just the difference between absorbed energy and
bond energy divided by 2 as we have 2 atoms after splitting:
E = (4.4 · 10^−19 − 4.0 · 10^−19)/2 = 0.2 · 10^−19 J = 2 * 10^-20 J
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

difference between absorbed energy and bond energy divided by 2
as we have 2 atoms after splitting:
E = (4.4 · 10^−19 − 4.0 · 10^−19)/2
= 0.2 · 10^−19 J
= 2 * 10^-20 J

Thanks and Regards
Hafsa Siddique
2 years ago
difference between absorbed energy and bond energy divided by 2
as we have 2 atoms after splitting:
E = (4.4 · 10^−19 − 4.0 · 10^−19)/2
= 0.2 · 10^−19 J
= 2 * 10^-20 J
 
Thanks and Regards